Assume a standard trade model. Which of the following statements is NOT true?

a) At the optimal output mix, the slope of the production possibility frontier equals the negative relative price.

b) If the economy produces more of one good, it has to produce less of the other good.

c) The production possibility frontier is convex.

d) The optimal output mix is realized where the isovalue line is tangent to the production possibility frontier.

Assume a firm faces the following cost function C(Q) = 10 · Q, where Q denotes output. Which of the following statements is true?

a) The average cost is decreasing in Q.

b) The marginal cost is constant.

c) There are economies of scale.

d) None of the previous statements is true.

Which two conditions have to be fulfilled to observe immiserizing growth?

a) The country has to be small and poor.

b) The country has to be large and the export sector has to grow much more than the import sector.

c) The country has to be poor and the growth has to be import biased.

d) The growth has to be either export biased or import biased

Assume a country produces tea and clothes. Both sectors use labor as an input in the production process and workers are mobile across sectors. In addition, the tea sector uses land as a specific factor, while the clothing industry uses capital as a specific factor. Assume there is a decrease in the price of tea. Which of the following statements is true?

a) Land owners are definitely better off as a result of the price change.

b) The real wage in terms of clothes increases as a result of the price change.

c) The real wage in terms of tea increases as a result of the price change.

d) None of the previous statements is true.

Assume a Ricardian model with two countries, France and the Netherlands, and two goods, wine and cheese. France has 60 units of labor available. France’s unit labor requirement to produce wine is 10. France’s unit labor requirement to produce cheese is 30. The Netherlands has 200 units of labor available. In the Netherlands the unit labor requirement to produce wine is 25. In the Netherlands the unit labor requirement to produce cheese is 5. Assume both countries can trade with each other. If the relative price of cheese in terms of wine is 4, how much wine will be produced in total (i.e., in both countries together)?

a) 0

b) 6

c) 8

d) 14
Answer #1

Assume a standard trade model. Which of the following statements is NOT true?

a) At the optimal output mix, the slope of the production possibility frontier equals the negative relative price.

b) If the economy produces more of one good, it has to produce less of the other good.

c) The production possibility frontier is convex.

d) The optimal output mix is realized where the isovalue line is tangent to the production possibility frontier.

reason: because the PPF is concave to the origin. Due to the increasing marginal cost of production.

Assume a firm faces the following cost function C(Q) = 10 · Q, where Q denotes output. Which of the following statements is true?

a) The average cost is decreasing in Q.

b) The marginal cost is constant.

c) There are economies of scale.

d) None of the previous statements is true.

because the derivative of the cost function is 10 which is constant.

Which two conditions have to be fulfilled to observe immiserizing growth?

a) The country has to be small and poor.

b) The country has to be large and the export sector has to grow much more than the import sector.

c) The country has to be poor and the growth has to be import biased.

d) The growth has to be either export biased or import biased

reason: this assumption is required as if the county is large and their export expands the price of the export falls too much and makes them worse off.

Assume a country produces tea and clothes. Both sectors use labor as an input in the production process and workers are mobile across sectors. In addition, the tea sector uses land as a specific factor, while the clothing industry uses capital as a specific factor. Assume there is a decrease in the price of tea. Which of the following statements is true?

a) Land owners are definitely better off as a result of the price change.

b) The real wage in terms of clothes increases as a result of the price change.

c) The real wage in terms of tea increases as a result of the price change.

d) None of the previous statements is true.

Reason: real wage in terms of tea is nominal wage/price of the tea hence the wage in terms of tea increase.

Assume a Ricardian model with two countries, France and the Netherlands, and two goods, wine and cheese. France has 60 units of labor available. France’s unit labor requirement to produce wine is 10. France’s unit labor requirement to produce cheese is 30. The Netherlands has 200 units of labor available. In the Netherlands the unit labor requirement to produce wine is 25. In the Netherlands the unit labor requirement to produce cheese is 5. Assume both countries can trade with each other. If the relative price of cheese in terms of wine is 4, how much wine will be produced in total (i.e., in both countries together)?

a) 0

b) 6

c) 8

d) 14

maximum amount of production possibitliy

France has 60 units of labour

they need 10 units to produce one unit of wine hence they can produce max 60/10=6 units of wine

they need 30 units to produce one unit of cheese so they produce 60/30=2 units of cheese

hence opportunity cost of producing one unit of cheese is 6/2=3 units of wine

Netherlands has 200 unit of labour

they need 25 units of labour to produce one unit of wine hence they can produce max 200/25=8 units of wine

they need 5 units of labour to produce one unit of cheese hence they can produce 200/5=40 units of cheese

hence relative price of cheese in netherlands=8/40=0.2 units of wine

hence the cost of producing cheese is less in Netherlands hence they will specialise in producing cheese and france will produce wine.

at price 4 wines per one unit of cheese, there will be no trade.

if relative price is 4wine per cheese then no trade will take place as it is a price which is beneficial for netherlads as they are in profit because their cost of producing one unit of cheese is only 0.2 units of wine. but this is not acceptable by france as the price is more than their own opportunity cost.

hence there will be no trade.

But maximum production of wine possible is 6+8=14